How to calculate cos x from its nth term in Maclaurin series??


1
Expert's answer
2021-11-29T05:26:11-0500

The Maclaurin series of "f(x)=\cos x" is

 

"f(x)=\sum_{n=0}^{\infty}(-1)^{n} \frac{x^{2 n}}{(2 n) !}"

 

Let us look at some details.

The Maclaurin series for f(x) in general can be found by

"f(x)=\sum_{n=0}^{\infty} \frac{f^{(n)}(0)}{n !} x^{n}"

Let us find the Maclaurin series for "f(x)=\cos x."

By taking the derivatives,

 "\begin{aligned}

&f(x)=\cos x \Rightarrow f(0)=\cos (0)=1 \\

&f^{\prime}(x)=-\sin x \Rightarrow f^{\prime}(0)=-\sin (0)=0 \\

&f^{\prime \prime}(x)=-\cos x \Rightarrow f^{\prime \prime}(0)=-\cos (0)=-1 \\

&f^{\prime \prime \prime}(x)=\sin x \Rightarrow f^{\prime \prime \prime}(0)=\sin (0)=0 \\

&f^{(4)}(x)=\cos x \Rightarrow f^{(4)}(0)=\cos (0)=1

\end{aligned}"

Since "f(x)=f^{(4)}(x)," the cycle of "\{1,0,-1,0\}" repeats itself.

So, we have the series

"f(x)=1-\frac{x^{2}}{2 !}+\frac{x^{4}}{4 !}-\cdots=\sum_{n=0}^{\infty}(-1)^{n} \frac{x^{2 n}}{(2 n) !}"